### v1

@@ 5.3 Synergies with electromagnetic and future gravitational-wave observatories
Bartel & Profumo (2024) have already shown that the detection of isolated MSPs via gravitational wave radiation will become tangible with third-generation ground-based detectors such as the Einstein Telescope. Looking ahead to space-based gravitational wave missions beyond LISA, several next-generation observatories have been proposed in response to European Space Agency’s call for future mission concepts. Among these, $\mu$-Ares (Sesana et al., 2021) and LISAmax (Martens et al., 2023) aim to improve sensitivity in the millihertz regime by at least an order of magnitude compared to LISA. Such a gain in sensitivity would significantly increase the number of detectable MSP binaries in the Galactic bulge and provide tighter constraints on their frequency evolution and sky localisation.
@@ 6 Conclusions
In this study, we explored whether gravitational-wave observations with LISA could contribute to the long-standing open question about the origin of the GeV $\gamma$-ray excess observed towards the Galactic bulge. Although the leading explanation in the literature points to dark matter annihilation, an alternative scenario involving a large population of MSPs has not been ruled out. The major challenge in testing the MSP hypothesis is that detecting pulsars in the inner Galaxy is notoriously difficult with electromagnetic observations, due to scattering and source confusion. We therefore asked whether the binary MSP population in the bulge could be probed through gravitational waves, independently of these limitations.
We showed that, at distances of 5–11 kpc (corresponding to the Galactic bulge), MSP binaries can be individually detected by LISA if they emit gravitational waves at frequencies above a few millihertz. This corresponds to orbital periods shorter than about 20 minutes. This regime is particularly interesting because it complements the capabilities of radio surveys, which are most sensitive to longer-period binaries and experience a sharp decline in detection efficiency for orbital periods below approximately one day. LISA therefore provides access to a region of parameter space that is currently inaccessible to electromagnetic observations.
To model the MSP binary population in the bulge, we considered two illustrative formation scenarios: one in which MSP binaries are deposited into the bulge through the accretion of disrupted globular clusters, and another in which they form in situ. The accreted scenario was constructed using a data-driven approach, with companion properties drawn from the well-established Tauris & Savonije (1999) relation and variations in the binary age distribution. In contrast, the in situ scenario was entirely theory-driven, based on binary population synthesis models that explore different channels of common-envelope evolution. While simplified, these scenarios span a plausible range of binary properties and enable us to examine how each might appear in LISA observations. In reality, the bulge MSP population likely includes contributions from both channels.
We find that both scenarios predict similar detection fractions, with the number of binaries detectable by LISA amounting to approximately $10^{-5}$$10^{-4}$ of the underlying MSP binary population. Although this fraction is small, even a single detection would imply the existence of a much larger unseen population, potentially numbering in the tens or hundreds of thousands. In the accreted scenario, MSP binaries typically have chirp masses of $\sim$0.4 M, while the in situ scenario produces systems with higher chirp masses of $\sim$0.9 M.
@@
The key challenge, however, lies in distinguishing MSP binaries from the far more numerous population of WDWD binaries. We illustrated this issue by overlaying a representative sample of WDWD binaries in Fig. 3. Although WDWD chirp masses typically peak at lower values, they remain abundant even in the regions of parameter space occupied by both accreted and in situ MSP binaries. As a result, while a small number of bulge MSP binaries are expected to be individually resolvable by LISA, distinguishing them from WDWDs will be the main obstacle in using gravitational-wave observations to test the MSP origin of the $\gamma$-ray excess. A promising strategy may be to use LISA detections as a targeted list for follow-up with the SKA, which could confirm the presence of MSPs in these binaries and provide critical evidence for or against their contribution to the $\gamma$-ray excess.
@@ Acknowledgements.
We thank Martyna Chruslinska, Mario Cadelano, Stephen Justham, and Vasily Belokurov for fruitful discussions and suggestions. AI thanks Anastasia Frantsuzova for answering multiple questions.

### v2

@@ 5.4 Synergies with electromagnetic and future gravitational-wave observatories
Bartel and Profumo (2024) have already shown that the detection of isolated MSPs via gravitational-wave radiation will become tangible with third-generation ground-based detectors such as the Einstein Telescope (Abac et al., 2025). Looking ahead to space-based gravitational-wave missions beyond LISA, several next-generation observatories have been proposed in response to European Space Agency’s call for future mission concepts. Among these, $\mu$-Ares (Sesana and others, 2021) and LISAmax (Martens et al., 2023) aim to improve sensitivity in the millihertz regime by at least an order of magnitude compared to LISA. Such a gain in sensitivity would significantly increase the number of detectable MSP binaries in the Galactic bulge and provide tighter constraints on their frequency evolution and sky localisation.
@@ 6 Conclusions
In this study, we have explored whether gravitational-wave observations with LISA could contribute to answering the long-standing open question about the origin of the gigaelectronvolt $\gamma$-ray excess observed towards the Galactic bulge. Although the leading explanation in the literature points to dark matter annihilation, an alternative scenario involving a large population of MSPs has not been ruled out. The major challenge in testing the MSP hypothesis is that detecting pulsars in the inner Galaxy is notoriously difficult with electromagnetic observations due to scattering and source confusion. We therefore investigated whether the binary MSP population in the bulge could be probed through gravitational waves with LISA independently of these limitations.
We showed that at distances of 5–11 kpc (corresponding to the Galactic bulge), MSP binaries can be individually detected by LISA if they emit gravitational waves at frequencies above a few millihertz. This corresponds to orbital periods shorter than about 20 minutes. This regime is particularly interesting because it complements the capabilities of radio surveys, which are most sensitive to longer-period binaries and experience a sharp decline in detection efficiency for orbital periods below approximately one day. LISA therefore provides access to a region of parameter space that is currently inaccessible to electromagnetic observations.
To model the MSP binary population in the bulge, we considered two illustrative formation scenarios: one in which MSP binaries are deposited into the bulge through the accretion of disrupted globular clusters and another in which they form in situ. The accreted scenario was constructed using a data-driven approach, with companion properties drawn from the well-established Tauris and Savonije (1999) relation and variations in the binary age distribution. In contrast, our in situ scenario is entirely theory driven, as it is based on binary population synthesis models that explore different channels of common-envelope evolution. While simplified, these scenarios span a plausible range of binary properties and enabled us to examine how each might appear in LISA observations. In reality, the bulge MSP population likely includes contributions from both channels.
We find that both scenarios predict similar detection fractions, with the number of binaries detectable by LISA amounting to approximately $10^{-5}$$10^{-4}$ of the underlying MSP binary bulge population. Although this fraction is small, even a single detection would imply the existence of a much larger unseen population, potentially numbering in the tens or hundreds of thousands. In the accreted scenario, MSP binaries typically have chirp masses of $\sim$0.4 M, while the in situ scenario produces systems with higher chirp masses of $\sim$0.9 M, though part of this contrast reflects our modelling assumptions.
@@
The key challenge, however, lies in distinguishing MSP binaries from the far more numerous population of WDWD binaries. We illustrated this issue by comparing a representative sample of WDWD binaries in Fig. 4. Although WDWD chirp masses typically peak at lower values, they remain abundant even in the regions of parameter space occupied by both accreted and in situ MSP binaries. As a result, while a small number of bulge MSP binaries are expected to be individually resolvable by LISA, distinguishing them from WDWDs will be the main obstacle in using gravitational-wave observations to test the MSP origin of the $\gamma$-ray excess. A promising strategy may be to use LISA detections as a targeted list for follow-up with the SKA, which could confirm the presence of MSPs in these binaries and provide critical evidence of, or against, their contribution to the $\gamma$-ray excess.